Suppose you have a pipe with a circumference of 8 cm and length of 20 cm and that 8 turns of a wire are wrapped around the pipe. What is the length of the wire?
The wire is wrapped around this cylinder in somewhat a diagonal from the top to the bottom. The wire does not cover the top or the bottom.
While working, I thought I should just multiply the 8 turns by the 8 cm circumference, but that does not give me the correct answer.

Answer by Alan3354(69443)

Suppose you have a pipe with a circumference of 8 cm and length of 20 cm and that 8 turns of a wire are wrapped around the pipe. What is the length of the wire?
The wire is wrapped around this cylinder in somewhat a diagonal from the top to the bottom. The wire does not cover the top or the bottom.
While working, I thought I should just multiply the 8 turns by the 8 cm circumference, but that does not give me the correct answer.
----------------------
The wire will form a helix. Each turn will advance along the pipe 1/8 of the length.
20/8 = 2.5 cm
Each turn will be similar to a right triangle with a base of 20 cm and a height of 2.5 cm. The hypotenuse will be sqrt(20^2 + 2.5^2) = sqrt(406.25)
--------------
*